A passenger train can travel 245 miles in the same amount of time it takes a freight train to travel 200 miles. If the rate of the passenger train is
15 MPH faster than the rate of the freight train, find the rate of each.
Set up a table and solve using an algebraic equation. Write write a complete sentence answer.
How do I even begin to start ?
Thank you

Answer by jorel555(1290)

Let r be the rate of the freight train. The passenger train goes r+15. So:
245/r+15=200/r
245r=200r+3000
45r=3000
r=66-2/3
r+15=81-2/3!!!!!!!!!!!!!!!!!!
245/r is 245 divided by r, which represents the time taken by the freight train......
Cross-multiplying 245/(r+15)=200/r, we get:
245r=200(r+15)
245r=200r+3000
etc..........